# Two-sample comparison through additive tree models for density ratios

### v2

@@ 4 Numerical Experiments
We conduct simulation experiments to evaluate the proposed methods—both tree boosting and Bayesian additive trees. We begin with scenarios involving bivariate distributions, and then consider higher-dimensional (20-dimensional) cases. We provide a comparison with several existing approaches. The first approach to be compared with is the “density-ratio trick” by inverting a binary classifier. In particular, we adopt AdaBoost as binary classification, as it uses the same class of additive tree ensembles to classify the two classes using essentially the exponential loss (Friedman et al., 2000). The difference between this approach and our boosting algorithms is essentially be due to the different loss functions employed. To apply the density-ratio trick, we first estimate the classification probabilities and then transform them into the density ratios by multiplying the sample size ratios. For implementation, we use the ada function provided in the R package ada (Culp et al., 2016). In addition to the “density-ratio trick”, we also compare our methods with two popular kernel-based density ratio estimators, namely KLEIP (Sugiyama et al., 2008) and uLSIF (Kanamori et al., 2009), implemented with the R package densratio (Makiyama, 2019).
For boosting, we assess both the FS and the GB algorithms. The learning rate $\nu$ is set to 0.01, and the maximum resolution of each tree is set to 4. The optimal number of trees is selected with the 5-fold cross-validation with the maximum number allowed set to 1000. For a fair comparison, AdaBoost is implemented under the same tuning parameters. For the generalized Bayesian additive tree model, we set the size of the ensemble to 200, following the standard implementation of the BART (Chipman et al., 2010). The MCMC sampler is initialized with trees containing only root nodes, and the size of the burn-in and the posterior sample is set to 2000 and 1000, respectively. In both types of the proposed models, we introduce 31 equally spaced cut points in each dimension. Further details on the implementation are provided in Supplementary Materials C.
@@
and for each setting we calculate the average MSE based on 50 simulated data sets.
4.1 2D simulations
@@ 1. Global shift
$X_{0}\sim N\left(\mu^{(0)},I_{2}\right),\ X_{1}\sim N\left(\mu^{(1)},I_{2}\right),$
where $\mu^{(0)}=(-0.5,-0.5)^{\prime}$ and $\mu^{(1)}=(0.5,0.5)^{\prime}$.
@@ 3. Local dispersion difference
A comparison of the MSEs is provided in Table 1, and the results indicate that our proposed methods (the boosting and the Bayesian additive trees) generally outperform the other approaches. We note that, for the performance of the classification approach with AdaBoost, the MSEs increase substantially when comparing the balanced and unbalanced sample-size settings. This confirms the common observation that inverting a binary classifier to attain the density ratio loses its efficacy when the sample sizes are unbalanced. By contrast, the proposed balancing loss remains robust even with unbalanced sample sizes.
@@
Table 1: A comparison of the MSEs and their standard errors in the two-dimensional scenarios. The methods are GB (gradient boosting), FS (forward stagewise), BAT (Bayesian additive trees), DRT (density-ratio trick based on AdaBoost), and the two kernel-based methods KLIEP and uLSIF.
@@
Figure 3: The estimated log-density ratios obtained from the algorithms considered in the two-dimensional examples for the local shift scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated pointwise are also displayed. The difference between FS and GB is minimal, so we present only the result for GB.
@@ 2. Dispersion Difference
Figure 4 shows the distribution of the first two latent factors, which account for the differences between the two groups, along with the corresponding true and estimated log density ratios. The latent factor model for the true densities represents a challenging case for density ratio estimation, particularly for tree-based methods that use axis-aligned partitions along the original dimensions of the sample space. Nevertheless, as shown in Table 2, the proposed methods still produce more accurate estimates, and the performance is robust to the unbalanced sample sizes. More importantly, in such challenging settings, uncertainty quantification becomes even more important to draw robust conclusions. The inferred density ratios must be gauged against appropriate measures of uncertainty to reach reliable inference. Our generalized Bayesian approach serves this need—the yellow/red regions in the lower quantiles and the blue points in the upper quantiles correspond to those falling outside the 95% credible bands, and these indeed align with the regions with true differences. Without this uncertainty quantification, the inferred different regions would have included much larger portions of the sample space across all of the estimates.
@@
Table 2: A comparison of the MSEs and their standard errors in the 20-dimensional scenarios. The methods are GB (gradient boosting), FS (forward stagewise), BAT (Bayesian additive trees), DRT (density-ratio trick using AdaBoost), and the two kernel-based methods KLIEP and uLSIF.
Figure 4: A comparison of the true log-density ratios and the estimated ratios projected onto the space of the latent variables $(z_{1},z_{2})$, for the location shift scenario ($n_{0}=n_{1}=5000$). The difference in the results for FS and GB is minimal, so we present only the result for GB.
@@ 5 Assessing the quality of synthesized microbiome compositions
In recent years, various generative models have been introduced to imitate the sampling mechanisms of various complex biomedical data. In particular, several approaches have been proposed for generating microbiome compositions (Sankaran et al., 2025). A common need is to assess the quality of the synthetic samples drawn from these models by comparing them with the actual observed data. We apply our two-sample comparison framework to synthetic samples simulated from different generative models and the corresponding observed samples, thereby assessing the quality of data generation in these models.
@@
The real microbiome data are obtained from the open-source Curated Metagenomic Data (Pasolli et al., 2017), available as the R package curatedMetagenomicData. We focus on the relative abundance data and apply the following subsetting rules: removing samples whose microbial relative abundance does not sum up to 1 (i.e., $\geq 0.99$); then, keep taxa with average relative abundance $\geq 0.1\%$; among those taxa, keep only samples that have zeros in fewer than 10% of taxa and whose mean abundance is at least the overall average. After subsetting, each row contains 123 columns of species-level relative abundance. We use one of the largest studies, the IBDMDB dataset (Lloyd-Price et al., 2019) in the database, as the data of interest. (The samples are in fact collected at multiple time points for each given subject, but the longitudinal nature of the data is important for scientific analysis but could be ignored for the purpose of this demonstration since we will only be focusing on assessing how well generative models can approximate the marginal distribution of the samples.) We split the data into training ($n=1166$) and testing ($n=292$) sets. The generative models are trained using the training data, and the testing data are used in the follow-up two-sample comparison with the synthetic samples. We include various popular generative models, including both classical parametric models and state-of-the-art deep neural network-based models. The parametric models considered are the Dirichlet model (Holmes et al., 2012; La Rosa et al., 2012) and the Dirichlet tree model (Dennis III, 1991; Wang and Zhao, 2017; Mao and Ma, 2022). The nonparametric generators based on neural networks include continuous normalizing flows via conditional flow matching (Lipman et al., 2022; Tong et al., 2023, ICFM) (code available at https://github.com/atong01/conditional-flow-matching), and a generative adversarial network model called MB-GAN (Rong et al., 2021) (code available at https://github.com/zhanxw/MB-GAN), which is tailored for microbiome data generation. Our objective here is not to argue which model is better than others, as such judgments are context-dependent and require more careful tuning of each model. Rather, we aim to simply provide an example that density ratio-based two-sample comparison can provide insights into how various generative models can be gauged given the observations that they generate.
Train v.s. Generated
Test v.s. Generated
@@
Figure 5: Principal Coordinates Analysis (PCoA) plots based on Bray-Curtis distance between samples. The blue and orange points represent the real sample and the generated sample, respectively. The first row presents the training sample vs the generated sample, and the second row presents the test sample vs the generated sample.
@@
We apply the generalized Bayesian additive trees and obtain a posterior sample of the density ratio evaluated at each of the real and the synthetic samples. We then compute the posterior mean along with a pointwise 95% credible band for the density ratio, as shown in Figure 6 (training sample vs synthetic sample) and Figure 7 (testing sample vs synthetic sample). The color indicates the estimated (log) density ratio. The estimated density ratios in both Figures 6 and 7 provide a clear contrast of the generation quality compared to the true sample. The estimated log-density ratio for the MB-GAN-generated sample versus the real sample is close to 0 across the support. The credible bands provided by our proposed method offer uncertainty quantification for the density ratio, demonstrating that 95% of the posterior mass of the log density ratio comfortably covers 0 over most of the sample space, which is not observed for the parametric models and ICFM.
Train v.s. Generated: Posterior mean of DRE
Train v.s. Generated: Lower 2.5% MCMC quantile of DRE
Train v.s. Generated: Upper 97.5% MCMC quantile of DRE
Test v.s. Generated: Posterior mean of DRE
Test v.s. Generated: Lower 2.5% MCMC quantile of DRE
Test v.s. Generated: Upper 97.5% MCMC quantile of DRE
@@ 6 Conclusion
We have introduced a new approach to learning the density ratio between two i.i.d. samples using an additive tree model as the function class. We showed that a new loss function called the balancing loss can be highly effective in learning the density ratio based on this function class. Not only does it provide effective boosting algorithms to attain functional (point) estimates, but it also provides a convenient means to generalized Bayesian inference on the density ratio, attaining uncertainty quantification which can be critical in challenging problems where the uncertainty in the learned density ratio can be substantial. The training algorithms we presented, both for boosting and for generalized Bayesian learning, are only minor modifications of their counterparts widely used in supervised learning based on this model class, and so their implementation would involve only minor adjustments to existing software for boosting and BART models.
Additive tree models do have their limitations. In particular, they use axis-aligned partitions and therefore can sometimes be less effective at capturing high-order interactions, and they generally do not perform well with very high-dimensional data, i.e., those involving more than a few hundred dimensions. In such scenarios, an effective strategy is to couple the tree-based approaches with strategies for learning effective representations of the data. Examples of such representations might include low-dimensional subspace assumptions (e.g., manifolds) or other highly non-linear embeddings through state-of-the-art neural network-based encoders. Even in such cases, we believe that learning the density ratio between two samples is a worthy objective and will typically be easier to achieve than learning the density of each of the two samples per se.
@@
We have implemented an R package BATTS for the proposed boosting and Bayesian additive tree algorithms, and it is available at https://github.com/MaStatLab/BATTS. R code for reproducing the numerical experiments is also provided at https://github.com/MaStatLab/BATTS_experiments.
@@ Acknowledgment
This research is funded in part by NIGMS grant R01-GM135440 and NSF grant DMS-2152999. Part of the research was completed when LM and YX were at Duke University. NA’s research is supported by the Japan Society for the Promotion of Science KAKENHI (grant numbers: 25K16618, 24H00142, and 25H00546). The authors thank Dr. Ganchao Wei for preprocessing the IBDMDB data and generating the synthetic samples.
Supplementary Materials
@@ Appendix A Proofs
A.1 Proof of Proposition 2.1
We first fix the finite tree partition $T_{k}$. After adding the piece-wise constant function $f_{k}$, the loss is given by
$l(w)=\sum_{A\in T_{k}}\left[e^{-\beta_{k}(A)}\int w^{-1}_{k-1}pd\mu+e^{\beta_{k}(A)}\int w_{k-1}qd\mu\right]$
@@
$e^{-\beta_{k}(A)}\int w^{-1}_{k-1}pd\mu=e^{\beta_{k}(A)}\int w_{k-1}qd\mu,$
and this condition yields the expression for the optimal $\beta_{k}(A)$. Furthermore, when the lower bound is attained for all $A\in T_{k}$, the loss can be expressed as
$l(w)=2\sqrt{\int w^{-1}_{k-1}pd\mu\int w_{k-1}qd\mu},$
@@
Also, we can derive the following corollary, by setting $T_{k}$ to the root tree.
Corollary A.1.

### v3

@@ 4 Numerical Experiments
We conduct simulation experiments to evaluate the proposed methods—both tree boosting and Bayesian additive trees. We begin with scenarios involving bivariate distributions, and then consider higher-dimensional (20-dimensional) cases, providing a comparison with several existing approaches. The first approach to be compared with is the “density-ratio trick” by inverting a binary classifier. In particular, we adopt AdaBoost for binary classification, as it uses the same class of additive tree ensembles to classify the two classes using essentially the exponential loss (Friedman et al., 2000). The difference between this approach and our boosting algorithms is essentially due to the different loss functions employed. To apply the density-ratio trick, we first estimate the classification probabilities and then transform them into the density ratios by multiplying the sample size ratios. For implementation, we use the ada function provided in the R package ada (Culp et al., 2016). In addition to the density-ratio trick, we evaluate the calibration method proposed in Cranmer et al. (2015) using AdaBoost as a discriminator. We also compare our methods with two popular kernel-based density ratio estimators, namely KLIEP (Sugiyama et al., 2008) and uLSIF (Kanamori et al., 2009), implemented with the R package densratio (Makiyama, 2019).
For boosting, we assess both the FS and the GB algorithms. The learning rate $\nu$ is set to 0.01, and the maximum resolution of each tree is set to 4. The optimal number of trees is selected with the 5-fold cross-validation with the maximum number allowed set to 1000. In both types of the boosting algorithms, we introduce 31 equally spaced cut points in each dimension. For a fair comparison, AdaBoost is implemented under the same tuning parameters. For the generalized Bayesian additive tree model, we set the size of the ensemble to 200, following the standard implementation of the BART (Chipman et al., 2010). The MCMC sampler is initialized with trees containing only root nodes, and the size of the burn-in and the posterior sample is set to 2000 and 1000, respectively. Further details on the implementation are provided in Supplementary Materials C.
@@
and for each setting we calculate the average of the error defined above based on 50 simulated data sets.
4.1 Two-dimensional Simulations
@@ 1. Global shift
$X_{0}\sim N\left(\mu^{(0)},I_{2}\right),\ X_{1}\sim N\left(\mu^{(1)},I_{2}\right)$, where $\mu^{(0)}=(-0.5,-0.5)^{\prime}$ and $\mu^{(1)}=(0.5,0.5)^{\prime}$.
@@ 3. Local dispersion difference
A comparison of the MSEs is provided in Table 1, and the results indicate that our proposed methods (the boosting and the Bayesian additive trees) generally show minimum errors. On the other hand, for the performance of the density ratio trick approach with AdaBoost, the MSEs increase substantially when comparing the balanced and unbalanced sample-size settings, reflecting the limitation discussed in Section 2.2. Also, the calibration method does not necessarily improve the performance, reflecting the fact that this approach involves a density estimation of the classification probabilities and thus can introduce an additional error. The two kernel-based methods (KLIEP and uLSIF) and the proposed methods based on the balancing loss are not strongly influenced by the unbalanced sample sizes, but the proposed method outperforms in all settings, due to the flexibility of the non-parametric models.
@@
Table 1: A comparison of the MSEs and their standard errors in the two-dimensional scenarios.
@@
uLSIF 1.068 0.422 0.383 0.410 0.142 0.211
(0.171) (0.096) (0.015) (0.022) (0.007) (0.010)
  • Notes: The methods are GB (gradient boosting), FS (forward stagewise), BAT (Bayesian additive trees), DRT and CDC (density-ratio trick and calibrated discriminative classifier based on AdaBoost), and the two kernel-based methods KLIEP and uLSIF.
@@
Figure 3: The estimated log-density ratios obtained from the algorithms considered in the two-dimensional examples for the local shift scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated pointwise are also displayed. The difference between FS and GB is minimal, so we present only the results for GB.
@@ 2. Dispersion Difference
Figure 4 shows the distribution of the first two latent factors in the first scenario, which account for the differences between the two groups, along with the corresponding true and estimated log-density ratios. The latent factor model for the true densities represents a challenging case for density ratio estimation, particularly for tree-based methods that use axis-aligned partitions along the original dimensions of the sample space. Nevertheless, as shown in Table 2, the proposed methods still produce more accurate estimates, and the performance is robust to the unbalanced sample sizes. More importantly, since estimating density ratios accurately becomes more challenging in such higher-dimensional settings, providing an uncertainty quantification for the estimates becomes even more crucial to conclude in which regions the two-sample difference exists. Our Bayesian approach serves this need by providing pointwise credible intervals for density ratios. The yellow/red regions in the lower quantiles and the blue regions in the upper quantiles correspond to the points at which the 95% credible bands exclude zero, indicating that the difference is significant at such points, and these indeed align with the regions with true differences.
Table 2: A comparison of the MSEs and their standard errors in the 20-dimensional scenarios.
@@
uLSIF 3.307 3.569 0.546 0.547
(0.070) (0.094) (0.001) (0.001)
  • Notes: The methods are GB (gradient boosting), FS (forward stagewise), BAT (Bayesian additive trees), DRT and CDC (density-ratio trick and calibrated discriminative classifier based on AdaBoost), and the two kernel-based methods KLIEP and uLSIF.
Figure 4: A comparison of the true log-density ratios and the estimated ratios projected onto the space of the latent variables $(z_{1},z_{2})$, for the location shift scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated pointwise are also displayed. The difference in the results for FS and GB is minimal, so we present only the results for GB.
@@ 5 Assessing the Quality of Synthesized Microbiome Compositions
In recent years, various generative models have been introduced to imitate sampling mechanisms of various complex biomedical data. In particular, several approaches have been proposed for generating microbiome compositions (Sankaran et al., 2025). A common need is to assess the quality of the synthetic samples drawn from these models by comparing them with the actual observed data. We apply our two-sample comparison framework to synthetic samples simulated from different generative models and the corresponding observed samples, thereby assessing the quality of data generation in these models.
@@
The real microbiome data are obtained from the open-source Curated Metagenomic Data (Pasolli et al., 2017), available as the R package curatedMetagenomicData. We focus on the relative abundance data and apply the following subsetting rules: remove observations whose microbial relative abundance does not sum up to 1 (i.e., $\geq 0.99$); then, keep taxa with average relative abundance $\geq 0.1\%$; among those taxa, keep only observations that have zeros in fewer than 10% of taxa and whose mean abundance is at least the overall average. After subsetting, each row contains 123 columns of species-level relative abundance. We use one of the largest studies, the IBDMDB dataset (Lloyd-Price et al., 2019) in the database, as the data of interest. (The samples are in fact collected at multiple time points for each given subject, and the longitudinal nature of the data is important for scientific analysis. However, we ignore this aspect for the purpose of this demonstration since we will only focus on assessing how well generative models can approximate the marginal distribution of the samples.) We split the data into training ($n=1166$) and testing ($n=292$) sets. The generative models are fitted using the training data, and the testing data are used in the follow-up two-sample comparison with the synthetic samples. We evaluate various popular generative models, including both classical parametric models and state-of-the-art deep neural network-based models. The parametric models considered are the Dirichlet model (Holmes et al., 2012; La Rosa et al., 2012) and the Dirichlet tree model (Dennis III, 1991; Wang and Zhao, 2017; Mao and Ma, 2022). The nonparametric generators based on neural networks include continuous normalizing flows via conditional flow matching (Lipman et al., 2022; Tong et al., 2023, ICFM) (code available at https://github.com/atong01/conditional-flow-matching), and a generative adversarial network model called MB-GAN (Rong et al., 2021) (code available at https://github.com/zhanxw/MB-GAN), which is tailored for microbiome data generation. Our objective here is not to argue which model is better than others, as such judgments are context-dependent and require more careful tuning of each model. Rather, we aim to simply provide an example that density ratio-based two-sample comparison can provide insights into how various generative models can be gauged given the observations that they generate.
Train vs. Generated
Test vs. Generated
@@
Figure 5: Principal Coordinates Analysis (PCoA) plots based on Bray-Curtis distance between observations. The blue and orange points represent the real sample and the generated sample, respectively. The first row presents the training sample vs. the generated sample, and the second row presents the test sample vs. the generated sample.
@@
We apply the generalized Bayesian additive trees and obtain a posterior sample of the density ratio evaluated at each pair of the real and the synthetic samples. We then compute the posterior mean along with the pointwise 95% credible band for the density ratio, as shown in Figure 6(a) (training sample vs synthetic sample) and Figure 7(a) (testing sample vs synthetic sample). The color indicates the estimated (log) density ratio. Figure 6(b) and Figure 7(b) provide the PCoA visualization for the observation-level credible interval coverage, specifically, whether the credible interval for each sample is below 0, covering 0, or above 0. Figure 6(c) and 7(c) provide the credible intervals for the top 20 observations with the largest absolute log-density ratio values. Supplementary Materials D provides detailed plots on the credible bands for all samples. Recall that a log-density ratio of zero indicates that the generated and real samples are equally represented at the distributional level, which is the defining property of an ideal generative model. The estimated density ratios in both Figures 6 and 7 provide a clear contrast of the generation quality compared to the true sample. Compared with other generative models, the estimated log-density ratio for the MB-GAN-generated sample versus the real sample is closer to 0 across the support, especially in the test data. As shown in Figure 7(b), the credible bands provided by our proposed method offer uncertainty quantification for the density ratio, demonstrating 95% pointwise credible intervals of the log-density ratio include zero over most of the sample space for the MB-GAN, which is not observed for the parametric models and ICFM. If we compare the credible bands in Figures 6 and 7, because the training data size ($n=1166$) is much larger than the testing data ($n=292$), the posterior credible bands are generally narrower in the training data (Figure 6). This is especially clear for MB-GAN, where there are more observations with narrow credible bands that do not cover 0, meaning the pointwise posterior mass is more concentrated and away from 0; whereas in the testing data, almost all MB-GAN observations have credible intervals that cover 0.
(a) Train vs. Generated: Posterior mean of DRE
(b) Train vs. Generated: CI is below 0 (blue), contains 0 (grey), or above 0 (red).
(c) Observation-level credible interval for top 20 subjects with the largest absolute value of log-density ratio. The y-axis shows the sample index.
Figure 6: For the training data: (Left) PCoA plots (x-axis and y-axis are the first and second principal components) colored by the density ratio estimates of the training sample density over the generated sample density, and whether the 2.5% and 97.5% credible bands cover 0; (Right) For a few selected observations with the largest absolute log-density ratio averaged over all methods, the point estimates and the credible intervals for each generative model.
(a) Test vs. Generated: Posterior mean of DRE
(b) Test vs. Generated: CI is below 0 (blue), contains 0 (grey), or above 0 (red).
(c) Observation-level credible interval for top 20 subjects with the largest absolute value of log-density ratio. The y-axis shows the sample index.
Figure 7: For the test data: (Left) PCoA plots (x-axis and y-axis are the first and second principal components) colored by the density ratio estimates of the testing sample density over the generated sample density, and whether the 2.5% and 97.5% credible bands cover 0; (Right) For a few selected observations with the largest absolute log-density ratio averaged over all methods, the point estimates and the credible intervals for each generative model.
@@ 6 Conclusion
We have introduced a new approach to learning the density ratio between two i.i.d. samples using the class of additive tree models and a new loss function called the balancing loss, which was shown to be highly effective in learning density ratios. Not only does it provide effective boosting algorithms to attain functional (point) estimates, but it also provides a convenient means for generalized Bayesian inference on the density ratio, attaining uncertainty quantification, which can be critical in challenging problems with limited sample size or high-dimensionality. The training algorithms we presented, both for boosting and for generalized Bayesian learning, can be introduced with modifications of their counterparts widely used in supervised learning based on this model class, and so their implementation would involve only minor adjustments to existing algorithms for boosting and BART models.
Additive tree models do have their limitations. In particular, they use axis-aligned partitions and therefore can sometimes be less effective at capturing high-order interactions, and they generally do not perform well with extremely high-dimensional data, i.e., those involving more than a few hundred dimensions. In such scenarios, an effective strategy is to couple the tree-based approaches with strategies for learning effective representations of the data. Examples of such representations might include low-dimensional subspace assumptions (e.g., manifolds) or other non-linear embeddings through state-of-the-art neural network-based encoders. Even in such cases, we believe that learning the density ratio between two samples is a worthy objective and will typically be easier to achieve than learning the density of each of the two samples per se.
@@
We have implemented an R package BATTS for the proposed boosting and Bayesian additive tree algorithms, and it is available at https://github.com/MaStatLab/BATTS .
@@ Acknowledgment
This research is funded in part by NIGMS grant R01-GM135440 and NSF grant DMS-2152999. Part of the research was completed when LM and YX were at Duke University. NA’s research is supported by the Japan Society for the Promotion of Science KAKENHI (grant numbers: 25K16618, 24H00142, and 25H00546). Portions of the manuscript were prepared with the assistance of ChatGPT (from GPT-4.1 to 5.4) for English-language editing.
Data Availability Statement
The real data, curated metagenomic data (Pasolli et al., 2017), used in Section 5 are openly available as an R package curatedMetagenomicData on GitHub and Bioconductor. The instruction and documentation are at https://waldronlab.io/curatedMetagenomicData/, DOI: 10.18129/B9.bioc.curatedMetagenomicData.
@@ Appendix A Proofs
A.1 Proof of Proposition 1
We first fix the finite tree partition $T_{k}$. After adding the piecewise constant function $f_{k}$, the loss is given by
$l(w)=\sum_{A\in T_{k}}\left[\exp(-\beta_{k}(A))\int_{A}w^{-1}_{k-1}p\,d\mu+\exp(\beta_{k}(A))\int_{A}w_{k-1}q\,d\mu\right]$
@@ A.1 Proof of Proposition 1
$\exp(-\beta_{k}(A))\int_{A}w^{-1}_{k-1}p\,d\mu=\exp(\beta_{k}(A))\int_{A}w_{k-1}q\,d\mu,$
and this condition yields the expression of the optimal $\beta_{k}(A)$. Furthermore, when the lower bound is attained for all $A\in T_{k}$, the loss can be expressed as
$l(w)=2\sum_{A\in T_{k}}\sqrt{\int_{A}w^{-1}_{k-1}p\,d\mu\int_{A}w_{k-1}q\,d\mu},$
@@
Also, we can derive the following corollary by considering the case in which the $T_{k}$ is the root tree.
Corollary 1.